Find the value for sin (x/2) if the following conditions hold: cos(x)= -1/3and x is in quadrant two.

Answer by stanbon(75887)

Find the value for sin (x/2) if the following conditions hold: cos(x)= -1/3and x is in quadrant two.
---------------------------
cos(x/2) = sqrt[(1+cos(x))/2] = sqrt[(2+(-1/3))/2] = sqrt[(5/6)]
===================
Cheers,
Stan H.
--------------